# Sum of Squares Decompositions for Structured Biquadratic Forms

### v1

@@ Abstract
This paper studies sum-of-squares (SOS) representations for structured biquadratic forms. We prove that diagonally dominated symmetric biquadratic tensors are always SOS. For the special case of symmetric biquadratic forms, we establish necessary and sufficient conditions for positive semi-definiteness of monic symmetric biquadratic forms, characterize the geometry of the corresponding psd cone, and prove that every such psd form is SOS. We also formulate conjectures regarding SOS representations for symmetric M-biquadratic tensors and symmetric B0-biquadratic tensors, discussing their likelihood and potential proof strategies. Our results advance the understanding of when positive semi-definiteness implies sum-of-squares decompositions for structured biquadratic forms.
Key words. Biquadratic forms, sum-of-squares, positive semi-definiteness, sos rank, bilinear forms.
@@ 1 Introduction
    We focus on the special case of symmetric biquadratic forms. Notably, Choi’s counterexample [3] is not symmetric. We establish necessary and sufficient conditions for positive semi-definiteness of monic symmetric biquadratic forms in the $3\times 3$ case (Theorem 5.2), characterize the geometry of the corresponding psd cone (Theorem 5.3), and prove that every $3\times 3$ symmetric psd biquadratic form is SOS (Theorem 5.5). We further conjecture that such forms can be expressed as the sum of at most five squares of bilinear forms (Conjecture 5.6).
The rest of this paper is organized as follows. Section 2 is devoted to diagonally dominated tensors, where we prove Theorem 2.1. Sections 3 and 4 introduce M- and $\mathrm{B}_{0}$-biquadratic tensors and present our conjectures. Section 5 formally defines $x$-symmetric and $y$-symmetric biquadratic forms and studies monic symmetric biquadratic forms. Section 6 offers further remarks and discusses directions for future research.
We denote $(1,\cdots,m)$ as $[m]$.
@@ 2 Diagonally Dominated Symmetric Biquadratic Tensors
Suppose that $\mathcal{A}=(a_{ijkl})$, where $a_{ijkl}\in\mathds{R}$ for $i,k\in[m],j,l\in[n]$. Then $\mathcal{A}$ is called an $m\times n$ biquadratic tensor. If
@@
for $i,k\in[m],j,l\in[n]$, then $\mathcal{A}$ is called a symmetric biquadratic tensor. Denote the set of all $m\times n$ biquadratic tensors by $BQ(m,n)$, and the set of all $m\times n$ symmetric biquadratic tensors by $SBQ(m,n)$. If the entries of $\mathcal{A}$ satisfy
@@
$M_{(i,j),(k,l)}=a_{ijkl},\qquad\text{for }i,k\in[m],\;j,l\in[n],$
@@
Now, define the vector $\mathbf{z}\in\mathds{R}^{mn}$ by $z_{(i,j)}=x_{i}y_{j}$ for $\mathbf{x}\in\mathds{R}^{m}$, $\mathbf{y}\in\mathds{R}^{n}$. Then the biquadratic form becomes:
$B(\mathbf{x},\mathbf{y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}x_{i}x_{k}y_{j}y_{l}=\mathbf{z}^{\top}M\mathbf{z}.$
@@
$M_{(i,j),(i,j)}\geq\sum_{(k,l)\neq(i,j)}|M_{(i,j),(k,l)}|\qquad\text{for all }(i,j)\in[m]\times[n],$
@@
A classical result states that any symmetric diagonally dominant matrix with nonnegative diagonal entries is positive semidefinite. Hence $M\succeq 0$, and thus $B(\mathbf{x},\mathbf{y})=\mathbf{z}^{\top}M\mathbf{z}\geq 0$ for all $\mathbf{x},\mathbf{y}$, confirming that $\mathcal{A}$ is psd.
@@
$M=\sum_{p=1}^{mn}\alpha_{p}\mathbf{e}_{p}\mathbf{e}_{p}^{\top}+\sum_{p<q}\beta_{pq}(\mathbf{e}_{p}+s_{pq}\mathbf{e}_{q})(\mathbf{e}_{p}+s_{pq}\mathbf{e}_{q})^{\top},$
@@
  • If $p=(i,j)$, then $\mathbf{e}_{p}\mathbf{e}_{p}^{\top}$ corresponds to $(x_{i}y_{j})^{2}$
  • If $p=(i,j)$ and $q=(k,l)$, then $(\mathbf{e}_{p}+s_{pq}\mathbf{e}_{q})(\mathbf{e}_{p}+s_{pq}\mathbf{e}_{q})^{\top}$ corresponds to $(x_{i}y_{j}+s_{pq}x_{k}y_{l})^{2}$
@@
$B(\mathbf{x},\mathbf{y})=\mathbf{z}^{\top}M\mathbf{z}=\sum_{r}({\bf w}_{r}^{\top}\mathbf{z})^{2},$
where each ${\bf w}_{r}^{\top}\mathbf{z}$ is a bilinear form in $\mathbf{x}$ and $\mathbf{y}$. This shows that $\mathcal{A}$ is an sos tensor. ∎
3 Symmetric M-Biquadratic Tensors
Let $\mathcal{A}=(a_{ijkl})\in SBQ(m,n)$, $\mathbf{x}\in\mathds{R}^{m}$ and $\mathbf{y}\in\mathds{R}^{n}$. Then $\mathcal{A}\cdot\mathbf{y}\mathbf{x}\mathbf{y})\in\mathds{R}^{m}$ and $(\mathcal{A}\cdot\mathbf{y}\mathbf{x}\mathbf{y}))_{i}=\sum_{k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}y_{j}x_{k}y_{l}$ for $i\in[m]$, $(\mathcal{A}\mathbf{x}\cdot\mathbf{x}\mathbf{y}))_{j}=\sum_{i,k=1}^{m}\sum_{l=1}^{n}a_{ijkl}x_{i}x_{k}y_{l}$ for $j\in[n]$. If there are $\lambda\in\mathds{R}$, $\mathbf{x}\in\mathds{R}^{m}$, $\|\mathbf{x}\|_{2}=1$, $\mathbf{y}\in\mathds{R}^{n}$, $\|\mathbf{y}\|_{2}=1$, such that
$\mathcal{A}\cdot\mathbf{y}\mathbf{x}\mathbf{y}=\lambda\mathbf{x},\ \mathcal{A}\mathbf{x}\cdot\mathbf{x}\mathbf{y}=\lambda\mathbf{y},$
then $\lambda$ is called an M-eigenvalue of $\mathcal{A}$, with $\mathbf{x}$ and $\mathbf{y}$ as its M-eigen vectors. M-eigenvalues of symmetric biquadratic tensors were introduced by Qi, Dai and Han [11] in 2009. It was proved there that a symmetric biquadratic tensor always has M-eigenvalues, and it is psd if and only if all of its M-eigenvalues are nonnegative.
@@ 3 Symmetric M-Biquadratic Tensors
A biquadratic tensor $\mathcal{A}$ in $BQ(m,n)$ is called a Z-biquadratic tensor if all of its off-diagonal entries are nonpositive. If $\mathcal{A}$ is a Z-biquadratic tensor, then it can be written as $\mathcal{A}=\alpha\mathcal{I}$-$\mathcal{B}$, where $\mathcal{I}$ is the M-identity tensor in $BQ(m,n)$, and $\mathcal{B}$ is a nonnegative biquadratic tensor. By [8], $\mathcal{B}$ has an M-eigenvalue. Denote ${\lambda_{\max}}(\mathcal{B})$ as the largest M eigenvalue of $\mathcal{B}$. If $\alpha\geq{\lambda_{\max}}(\mathcal{B})$, then $\mathcal{A}$ is called an M-biquadratic tensor. It was proved in [8], a symmetric biquadratic M-tensor is psd.
@@
4 Symmetric B0-Biquadratic Tensors
@@
For more discussion on biquadratic tensors, see [5, 13, 14]
@@ 5 Symmetric Biquadratic Forms
Suppose that $\mathcal{A}=(a_{ijkl})$, where $a_{ijkl}\in\mathds{R}$ for $i,k\in[m],j,l\in[n]$, is an $m\times n$ biquadratic tensor. Let
$P(\mathbf{x},\mathbf{y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}x_{i}y_{j}x_{k}y_{l},$ (5)
where $\mathbf{x}\in\mathds{R}^{m}$ and $\mathbf{y}\in\mathds{R}^{n}$. Then $P$ is called a biquadratic form. If $P(\mathbf{x},\mathbf{y})\geq 0$ for all $\mathbf{x}\in\mathds{R}^{m}$ and $\mathbf{y}\in\mathds{R}^{n}$, then $P$ is called positive semi-definite (psd). If
$P(\mathbf{x},\mathbf{y})=\sum_{p=1}^{r}f_{p}(\mathbf{x},\mathbf{y})^{2},$
@@ 5.1 Historical Context and Symmetric Forms
Suppose $P(\mathbf{x},\mathbf{y})=P(x_{1},\cdots,x_{m},y_{1},\cdots,y_{n})$. If
$P(x_{1},\cdots,x_{m},y_{1},\cdots,y_{n})=P(x_{\sigma(1)},\cdots,x_{\sigma(m)},y_{1},\cdots,y_{n})$
@@
We call such a form a monic symmetric biquadratic form. Note that if the diagonal coefficients are negative, then $P$ is not psd. If the diagonal coefficients are zero, then $P$ is not psd unless $a=b=c=0$. This exceptional cases are excluded from our analysis, as this paper focuses on psd and sos forms.
@@ 5.2 Positive Semi-definiteness of Monic Symmetric Forms
Let $\mathcal{A}\in{SBQ(m,n)}$ be the biquadratic tensor corresponding to $P$. Then $\lambda$ is called an M-eigenvalue of $\mathcal{A}$, with $\mathbf{x}$ and $\mathbf{y}$ as its M-eigenvectors if $\|\mathbf{x}\|_{2}=1$, $\|\mathbf{y}\|_{2}=1$,
$\beta_{x}\mathbf{1}_{m}+(1-a+(b-c)(\alpha_{y}^{2}-1))\mathbf{x}=\lambda\mathbf{x},$
and
$\beta_{y}\mathbf{1}_{n}+(1-b+(a-c)(\alpha_{x}^{2}-1))\mathbf{y}=\lambda\mathbf{y}.$
Here, $\alpha_{x}=\mathbf{1}_{m}^{\top}\mathbf{x}$, $\alpha_{y}=\mathbf{1}_{n}^{\top}\mathbf{y}$, $\beta_{x}=\alpha_{x}(a+c(\alpha_{y}^{2}-1))$, $\beta_{y}=\alpha_{y}(b+c(\alpha_{x}^{2}-1))$, $\mathbf{1}_{m}$ and $\mathbf{1}_{n}$ denote the all ones vectors of length $m$ and $n$, respectively. Therefore, $\mathbf{x}$ is an M-eigenvector of $\mathcal{A}$ only if $\beta_{x}=0$ or $\mathbf{x}=\frac{1}{\sqrt{m}}\mathbf{1}_{m}$. Similarly, $\mathbf{y}$ is an M-eigenvector of $\mathcal{A}$ only if $\beta_{y}=0$ or $\mathbf{y}=\frac{1}{\sqrt{n}}\mathbf{1}_{n}$. In the following, we prove conditions (i)–(iv) by enumerating all M-eigenvectors.
For condition (i), take $\mathbf{x}=\frac{1}{\sqrt{m}}\mathbf{1}_{m}$ and let $\mathbf{y}$ satisfy $\beta_{y}=0$. Then we have $\lambda=1-b+(m-1)(a-c)$ is an M-eigenvalue of $\mathcal{A}$. Since $P$ is psd, it follows that $1-b+(m-1)(a-c)\geq 0$.
Condition (ii) follows by taking $\mathbf{y}=\frac{1}{\sqrt{n}}\mathbf{1}_{n}$ and letting $\mathbf{x}$ satisfy $\beta_{x}=0$. This leads to the M-eigenvalue $1-a+(n-1)(b-c)\geq 0$.
For condition (iii), let $\mathbf{x}$ satisfy $\beta_{x}=0$ and $\mathbf{y}$ satisfy $\beta_{y}=0$, respectively. A direct computation gives $\lambda=1-a-b+c\geq 0$ and $\lambda=1-\frac{ab}{c}\geq 0$ when $c\neq 0$, $\frac{a}{c}\in[1-n,1]$, and $\frac{b}{c}\in[1-m,1]$.
For condition (iv), let $\mathbf{x}=\frac{1}{\sqrt{m}}\mathbf{1}_{m}$ and $\mathbf{y}=\frac{1}{\sqrt{n}}\mathbf{1}_{n}$. A direct computation gives $\lambda=1+(m-1)a+(n-1)b+(m-1)(n-1)c\geq 0$.
Conditions (i)-(iv) have enumerated all M-eigenvalues. Based on the equivalence between positive semidefiniteness and the nonnegativity of the smallest M-eigenvalue, the conclusion of this theorem is obtained. This completes the proof. ∎
5.3 Geometry of the $3\times 3$ PSD Cone

### v2

@@ Abstract
This paper studies sum-of-squares (SOS) representations for structured biquadratic forms. We prove that diagonally dominated symmetric biquadratic tensors are always SOS. For the special case of symmetric biquadratic forms, we establish necessary and sufficient conditions for positive semi-definiteness of monic symmetric biquadratic forms, characterize the geometry of the corresponding PSD cone as a convex polyhedron, and prove that every such PSD form is SOS for any dimensions $m$ and $n$. We also formulate conjectures regarding SOS representations for symmetric M-biquadratic tensors and symmetric $\mathrm{B}_{0}$-biquadratic tensors, discussing their likelihood and potential proof strategies. Our results advance the understanding of when positive semi-definiteness implies sum-of-squares decompositions for structured biquadratic forms.
Keywords. Biquadratic forms, sum-of-squares, positive semi-definiteness, M-eigenvalues, symmetric bilinear forms.
@@ 1 Introduction
    We focus on the special case of symmetric biquadratic forms. Notably, Choi’s counterexample [2] is not symmetric. We establish necessary and sufficient conditions for positive semi-definiteness of monic symmetric biquadratic forms (Theorem 4.2), characterize the geometry of the corresponding PSD cone as a convex polyhedron (Theorem 4.3), and prove that every symmetric PSD biquadratic form is SOS for any dimensions $m$ and $n$ (Theorem 4.5). We conjecture that $3\times 3$ symmetric forms can be expressed as the sum of at most five squares of bilinear forms (Conjecture 4.7).
The rest of this paper is organized as follows. Section 2 is devoted to diagonally dominated tensors, where we prove Theorem 2.1. Section 3 presents conjectures on the SOS property for symmetric M-biquadratic tensors and symmetric $\mathrm{B}_{0}$-biquadratic tensors. Section 4 formally defines $x$-symmetric and $y$-symmetric biquadratic forms and studies monic symmetric biquadratic forms, including the general SOS result. Section 5 offers further remarks and discusses directions for future research.
We denote $\{1,\ldots,m\}$ as $[m]$.
@@ 2 Diagonally Dominated Symmetric Biquadratic Tensors
Suppose that $\mathcal{A}=(a_{ijkl})$, where $a_{ijkl}\in\mathbb{R}$ for $i,k\in[m],j,l\in[n]$. Then $\mathcal{A}$ is called an $m\times n$ biquadratic tensor. If
@@
for $i,k\in[m],j,l\in[n]$, then $\mathcal{A}$ is called a symmetric biquadratic tensor. Denote the set of all $m\times n$ biquadratic tensors by $BQ(m,n)$, and the set of all $m\times n$ symmetric biquadratic tensors by $SBQ(m,n)$.
If the entries of $\mathcal{A}$ satisfy
@@
$M_{(i,j),(k,l)}=a_{ijkl},\quad\text{for }i,k\in[m],\;j,l\in[n],$
@@
Now, define the vector ${\bf z}\in\mathbb{R}^{mn}$ by $z_{(i,j)}=x_{i}y_{j}$ for ${\bf x}\in\mathbb{R}^{m}$, ${\bf y}\in\mathbb{R}^{n}$. Then the biquadratic form becomes:
$B({\bf x},{\bf y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}x_{i}x_{k}y_{j}y_{l}={\bf z}^{\top}M{\bf z}.$
@@
$M_{(i,j),(i,j)}\geq\sum_{(k,l)\neq(i,j)}|M_{(i,j),(k,l)}|\quad\text{for all }(i,j)\in[m]\times[n],$
@@
A classical result states that any symmetric diagonally dominant matrix with nonnegative diagonal entries is positive semidefinite. Hence $M\succeq 0$, and thus $B({\bf x},{\bf y})={\bf z}^{\top}M{\bf z}\geq 0$ for all ${\bf x},{\bf y}$, confirming that $\mathcal{A}$ is PSD.
@@
$M=\sum_{p=1}^{mn}\alpha_{p}{\bf e}_{p}{\bf e}_{p}^{\top}+\sum_{p<q}\beta_{pq}({\bf e}_{p}+s_{pq}{\bf e}_{q})({\bf e}_{p}+s_{pq}{\bf e}_{q})^{\top},$
@@
  • If $p=(i,j)$, then ${\bf e}_{p}{\bf e}_{p}^{\top}$ corresponds to $(x_{i}y_{j})^{2}$
  • If $p=(i,j)$ and $q=(k,l)$, then $({\bf e}_{p}+s_{pq}{\bf e}_{q})({\bf e}_{p}+s_{pq}{\bf e}_{q})^{\top}$ corresponds to $(x_{i}y_{j}+s_{pq}x_{k}y_{l})^{2}$
@@
$B({\bf x},{\bf y})={\bf z}^{\top}M{\bf z}=\sum_{r}({\bf w}_{r}^{\top}{\bf z})^{2},$
where each ${\bf w}_{r}^{\top}{\bf z}$ is a bilinear form in $\bf x$} and $\bf y$}. This shows that $\mathcal{A}$ is an SOS tensor. ∎
3 Conjectures for Structured Biquadratic Tensors
This section presents two conjectures regarding the SOS property for two important classes of structured biquadratic tensors: symmetric M-biquadratic tensors and symmetric $\mathrm{B}_{0}$-biquadratic tensors.
3.1 Symmetric M-Biquadratic Tensors
Let $\mathcal{A}=(a_{ijkl})\in SBQ(m,n)$, ${\bf x}\in\mathbb{R}^{m}$ and ${\bf y}\in\mathbb{R}^{n}$. Then $\mathcal{A}\cdot{\bf y}{\bf x}{\bf y})\in\mathbb{R}^{m}$ and $(\mathcal{A}\cdot{\bf y}{\bf x}{\bf y}))_{i}=\sum_{k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}y_{j}x_{k}y_{l}$ for $i\in[m]$, $(\mathcal{A}{\bf x}\cdot{\bf x}{\bf y}))_{j}=\sum_{i,k=1}^{m}\sum_{l=1}^{n}a_{ijkl}x_{i}x_{k}y_{l}$ for $j\in[n]$. If there are $\lambda\in\mathbb{R}$, ${\bf x}\in\mathbb{R}^{m}$, $\|{\bf x}\|_{2}=1$, ${\bf y}\in\mathbb{R}^{n}$, $\|{\bf y}\|_{2}=1$, such that
$\mathcal{A}\cdot{\bf y}{\bf x}{\bf y}=\lambda{\bf x},\quad\mathcal{A}{\bf x}\cdot{\bf x}{\bf y}=\lambda{\bf y},$
then $\lambda$ is called an M-eigenvalue of $\mathcal{A}$, with $\bf x$} and $\bf y$} as its M-eigenvectors. M-eigenvalues of symmetric biquadratic tensors were introduced by Qi, Dai and Han [10] in 2009. It was proved there that a symmetric biquadratic tensor always has M-eigenvalues, and it is PSD if and only if all of its M-eigenvalues are nonnegative.
@@ 3.1 Symmetric M-Biquadratic Tensors
A biquadratic tensor $\mathcal{A}$ in $BQ(m,n)$ is called a Z-biquadratic tensor if all of its off-diagonal entries are nonpositive. If $\mathcal{A}$ is a Z-biquadratic tensor, then it can be written as $\mathcal{A}=\alpha\mathcal{I}$-$\mathcal{B}$, where $\mathcal{I}$ is the M-identity tensor in $BQ(m,n)$, and $\mathcal{B}$ is a nonnegative biquadratic tensor. By [7], $\mathcal{B}$ has an M-eigenvalue. Denote ${\lambda_{\max}}(\mathcal{B})$ as the largest M eigenvalue of $\mathcal{B}$. If $\alpha\geq{\lambda_{\max}}(\mathcal{B})$, then $\mathcal{A}$ is called an M-biquadratic tensor. It was proved in [7] that a symmetric biquadratic M-tensor is PSD.
@@
3.2 Symmetric $\mathrm{B}_{0}$-Biquadratic Tensors
@@
For more discussion on biquadratic tensors, see [4, 12, 13].
@@ 4 Symmetric Biquadratic Forms
Suppose that $\mathcal{A}=(a_{ijkl})$, where $a_{ijkl}\in\mathbb{R}$ for $i,k\in[m],j,l\in[n]$, is an $m\times n$ biquadratic tensor. Let
$P({\bf x},{\bf y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}x_{i}y_{j}x_{k}y_{l},$ (5)
where ${\bf x}\in\mathbb{R}^{m}$ and ${\bf y}\in\mathbb{R}^{n}$. Then $P$ is called a biquadratic form. If $P({\bf x},{\bf y})\geq 0$ for all ${\bf x}\in\mathbb{R}^{m}$ and ${\bf y}\in\mathbb{R}^{n}$, then $P$ is called positive semi-definite (PSD). If
$P({\bf x},{\bf y})=\sum_{p=1}^{r}f_{p}({\bf x},{\bf y})^{2},$
@@ 4.1 Historical Context and Symmetric Forms
Suppose $P({\bf x},{\bf y})=P(x_{1},\ldots,x_{m},y_{1},\ldots,y_{n})$. If
$P(x_{1},\ldots,x_{m},y_{1},\ldots,y_{n})=P(x_{\sigma(1)},\ldots,x_{\sigma(m)},y_{1},\ldots,y_{n})$
@@
We call such a form a monic symmetric biquadratic form.
Note that if the diagonal coefficients are negative, then $P$ is not PSD. If the diagonal coefficients are zero, then $P$ is not PSD unless $a=b=c=0$. These exceptional cases are excluded from our analysis, as this paper focuses on PSD and SOS forms.
@@ 4.2 Positive Semi-definiteness of Monic Symmetric Forms
Let $\mathcal{A}\in{SBQ(m,n)}$ be the biquadratic tensor corresponding to $P$. Then $\lambda$ is called an M-eigenvalue of $\mathcal{A}$, with $\bf x$} and $\bf y$} as its M-eigenvectors if $\|{\bf x}\|_{2}=1$, $\|{\bf y}\|_{2}=1$, and
$\beta_{x}\mathbf{1}_{m}+(1-a+(b-c)(\alpha_{y}^{2}-1)){\bf x}$ $=\lambda{\bf x},$ (7)
$\beta_{y}\mathbf{1}_{n}+(1-b+(a-c)(\alpha_{x}^{2}-1)){\bf y}$ $=\lambda{\bf y}.$ (8)
Here, $\alpha_{x}=\mathbf{1}_{m}^{\top}{\bf x}$, $\alpha_{y}=\mathbf{1}_{n}^{\top}{\bf y}$, $\beta_{x}=\alpha_{x}(a+c(\alpha_{y}^{2}-1))$, $\beta_{y}=\alpha_{y}(b+c(\alpha_{x}^{2}-1))$, $\mathbf{1}_{m}$ and $\mathbf{1}_{n}$ denote the all ones vectors of length $m$ and $n$, respectively. Therefore, $\bf x$} is an M-eigenvector of $\mathcal{A}$ only if $\beta_{x}=0$ or ${\bf x}=\frac{1}{\sqrt{m}}\mathbf{1}_{m}$. Similarly, $\bf y$} is an M-eigenvector of $\mathcal{A}$ only if $\beta_{y}=0$ or ${\bf y}=\frac{1}{\sqrt{n}}\mathbf{1}_{n}$.
In the following, we prove conditions (i)–(v) by enumerating all M-eigenvectors.
For condition (i), take ${\bf x}=\frac{1}{\sqrt{m}}\mathbf{1}_{m}$ and choose $\bf y$} such that $\beta_{y}=0$. Then it follows from (8) that $\lambda=1-b+(m-1)(a-c)$ is an M-eigenvalue of $\mathcal{A}$. Indeed, substituting $\lambda=1-b+(m-1)(a-c)$ to (7) yields $\alpha_{y}=0$. Since $P$ is PSD, it follows that $1-b+(m-1)(a-c)\geq 0$.
Condition (ii) follows by taking ${\bf y}=\frac{1}{\sqrt{n}}\mathbf{1}_{n}$ and letting $\bf x$} satisfy $\beta_{x}=0$. This leads to the M-eigenvalue $1-a+(n-1)(b-c)\geq 0$.
For conditions (iii) and (v), let $\bf x$} satisfy $\beta_{x}=0$ and $\bf y$} satisfy $\beta_{y}=0$, respectively. Consider the following subcases:
  • (iii.1)
    $\alpha_{x}=0$ and $\alpha_{y}=0$. Then we may derive that $\lambda=1-a-b+c$.
  • (iii.2)
    $\alpha_{x}=0$ and $\alpha_{y}\neq 0$. Then it follows from (8) that $\lambda=1-a-b+c$.
  • (iii.3)
    $\alpha_{x}\neq 0$ and $\alpha_{y}=0$. Then it follows from (7) that $\lambda=1-a-b+c$.
  • (iii.4)
    $\alpha_{x}\neq 0$, $\alpha_{y}\neq 0$, and $c=0$. Then it follows from $\beta_{x}=0$ and $\beta_{y}=0$ that $a=b=0$. Consequently, there is $\lambda=1-a-b+c$.
  • (iii.5)
    $\alpha_{x}\neq 0$, $\alpha_{y}\neq 0$, and $c\neq 0$. It follows from $\beta_{x}=0$ and $\beta_{y}=0$ that $\alpha_{x}^{2}=\frac{c-b}{c}$ and $\alpha_{y}^{2}=\frac{c-a}{c}$. This further implies $\lambda=1-\frac{ab}{c}\geq 0$. Furthermore, by Cauchy inequality and $\|{\bf x}\|=\|{\bf y}\|=1$, we have $\alpha_{x}^{2}\in(0,m],\alpha_{y}^{2}\in(0,n]$. Therefore, $\frac{a}{c}\in[1-n,1)$, and $\frac{b}{c}\in[1-m,1)$.
For condition (iv), let ${\bf x}=\frac{1}{\sqrt{m}}\mathbf{1}_{m}$ and ${\bf y}=\frac{1}{\sqrt{n}}\mathbf{1}_{n}$. A direct computation gives $\lambda=1+(m-1)a+(n-1)b+(m-1)(n-1)c\geq 0$.
Conditions (i)–(v) have enumerated all M-eigenvalues. Based on the equivalence between positive semidefiniteness and the nonnegativity of the smallest M-eigenvalue, the conclusion of this theorem is obtained. This completes the proof. ∎
4.3 The PSD Cone is a Convex Polyhedron
Although Theorem 4.2 contains a nonlinear condition (v), it turns out that for any $m,n\geq 2$, the set of coefficients $(a,b,c)$ for which $P$ is PSD is actually a convex polyhedron defined only by the four linear inequalities (i)–(iv). The vertices of this polyhedron are easily computed.